# A Hybrid Systems Model of Feedback Optimization for Linear Systems: Convergence and Robustness

### v4

@@ VI Simulation Results
Figure 2: The left-hand plot shows the convergence of the state $x$ under the dynamics $\mathcal{H}_{FO}$ from the initial condition in (90). We see that $x$ attains an asymptotic error of $5\cdot 10^{-9}$, which illustrates that it converges quite close to its desired value. The right-hand plot shows the convergence of the state $x$ under the perturbed dynamics $\mathcal{H}_{FO}^{\rho}$ from the initial condition in (90). In this case, $x$ attains an asymptotic error of $1.52$, even though there are simultaneously perturbations to the timer dynamics and the model of the underlying LTI system.
In this section we present two sets of simulation results: one with the nominal system $\mathcal{H}_{FO}$ and one with the perturbed system $\mathcal{H}_{FO}^{\rho}$. We consider the LTI dynamics
@@
subject to $y=Hu+d,\,\,u\in\mathcal{U},\,\,y\in\mathbb{R}^{4},$
where we use $Q_{u}=0.08$, $Q_{y}=0.3I_{4}$, $\mathcal{U}=[0,12]$, $d=(0.2,0.2,0.2,0.2)^{\top}$, and $\hat{y}=(4,4,4,4)^{\top}$.
For simulations, the Hybrid Equations Toolbox (Version 3.0.0.76) was used, along with the initial conditions
$x(0,0)=(0,3,10,15)^{\top},u(0,0)=0,\\ y_{s}(0,0)=(0.2,3.2,10.2,15.2)^{\top},z(0,0)=0,\\ \tau_{c}(0,0)=0.175,\penalty 10000\ \tau_{g}(0,0)=0.05,$ (90)
where $\tau_{g,comp}=0.05,\tau_{c,\min}=0.15$, and $\tau_{c,\max}=0.20$ with stepsize $\gamma=0.35\leq\frac{2}{\lambda_{min}(Q_{u})+L}$. We see in the left-hand plot of Figure 2 that $\|x-\tilde{x}\|$ converges to $5\cdot 10^{-9}$, and thus $x$ is asymptotically close to its desired value.
We next consider the perturbed case. Let $J_{a,b}$ denote the matrix of all ones in $\mathbb{R}^{a\times b}$. For the timers, we use perturbations of the form $\theta_{c,\min}=\theta_{c,\max}=\theta_{g,comp}=\kappa_{c}=\kappa_{g}=\chi$ for several values of $\chi$, and for the LTI dynamics we use perturbations of the form $\hat{A}=\mu J_{n\times n}$, $\hat{B}=\mu J_{n\times m}$, and $\hat{H}=\mu J_{p\times m}$ for several values of $\mu$. The values of $\chi$ and $\mu$ used in simulations are shown in Table I, along with the values they induce in the steady-state error $\|x-\tilde{x}\|$. Using these values, the perturbation $\rho$ from (84) can be quite large because it is the maximum of several terms, including $\|\hat{A}x\|$, $\|\hat{B}u\|$, and $\|\hat{H}u\|$, which can be large because they depend on $x$ or $u$.
TABLE I: Values of perturbations and resulting asymptotic error for the perturbed system $\mathcal{H}_{FO}^{\rho}$.
Limiting value of $\|x-\tilde{x}\|$
$\chi=0.2$, $\mu=0.01$ 0.28
$\chi=0.4$, $\mu=0.02$ 0.51
$\chi=0.6$, $\mu=0.03$ 0.21
$\chi=0.8$, $\mu=0.04$ 0.64
$\chi=0.95$, $\mu=0.05$ 0.80
Across all runs, $\|x-\tilde{x}\|$ ranged from $0.28$ to $0.80$, which indicates that $\mathcal{H}_{FO}^{\rho}$ successfully reduces error and rejects disturbances, even under perturbations $\rho$ whose size is comparable to the size of other terms in the dynamics.

### v5

@@ VI Simulation Results
Figure 2: The left-hand plot shows the convergence of the state $x$ under the dynamics $\mathcal{H}_{FO}$ from the initial condition in (85). We see that $x$ attains an asymptotic error of $3.26\cdot 10^{-15}$, which illustrates that it converges quite close to its desired value. The right-hand plot shows the convergence of the state $x$ under the perturbed dynamics $\mathcal{H}_{FO}^{\iota}$ from the initial condition in (85). In this case, $x$ attains an asymptotic error of at most $12.22$ relative to the errors from $\mathcal{H}_{FO}$ under large model and timer perturbations. under these perturbations, Assumption 1 still guarantees exponential convergence.
In this section we present two sets of simulation results: one with the nominal system $\mathcal{H}_{FO}$ and one with the perturbed system $\mathcal{H}_{FO}^{\iota}$. We consider the LTI dynamics
@@
[…]
where we use $Q_{u}=0.08$, $Q_{y}=0.3I_{4}$, $\mathcal{U}=[-5,5]$, $d=(0.2,0.2,0.2,0.2)^{\top}$, and $\hat{y}=(4,4,4,4)^{\top}$.
For simulations, the Hybrid Equations Toolbox (Version 3.0.0.76 [25] ) was used, along with the initial conditions
$x(0,0)=(0,5,10,15)^{\top},u(0,0)=0,\\ y_{s}(0,0)=(0.2,5.2,10.2,15.2)^{\top},z(0,0)=0,\\ \tau_{c}(0,0)=0.175,\penalty 10000\ \tau_{g}(0,0)=0.05,$ (85)
where $\tau_{g,comp}=0.05,\tau_{c,\min}=0.15$, and $\tau_{c,\max}=0.20$ which ensures at least three iterations with stepsize $\gamma=0.35\leq\frac{2}{\lambda_{min}(Q_{u})+L}$. We see in the left-hand plot of Figure 2 that $\|x-\tilde{x}\|$ converges to $3.26\cdot 10^{-15}$, and thus $x$ is asymptotically close to its desired value.
We next consider the perturbed case. Let $J_{a,b}$ denote the matrix of all ones in $\mathbb{R}^{a\times b}$. For the timers, we use perturbations of the form $\theta_{c,\min}=\theta_{c,\max}=\theta_{g,comp}=0.5$ and $\kappa_{c}=\kappa_{g}=\kappa$ for several values of $\kappa$, and for the LTI dynamics we use perturbations of the form $\hat{A}=\mu J_{n\times n}$, $\hat{B}=\mu J_{n\times m}$, and $\hat{H}=\mu J_{p\times m}$ for several values of $\mu$. The values of $\kappa$ and $\mu$ used in simulations are shown in Table I, along with the values they induce in the steady-state error $\|x-\tilde{x}\|$ (from $0.86$ to $12.22$). Using these values, the perturbation $\rho$ from (79) can be quite large because it is the maximum of several terms, including $\|\hat{A}x\|$, $\|\hat{B}u\|$, and $\|\hat{H}u\|$, which can be large because they depend on $x$ or $u$. A second-order regression takes the form $\|\mathbf{x}-\tilde{\mathbf{x}}\|\approx 0.217+6.054\mu+8.036\mu^{2}$, showing that the steady-state error is dependent on the model perturbations rather than the time errors. This agrees with the right-hand plot of Figure 2 where $\kappa$ affected the convergence rate.
TABLE I: Values of perturbations and the asymptotic error for the perturbed system $\mathcal{H}_{FO}^{\iota}$ relative to $\mathcal{H}_{FO}$.
Value of $\|x-\tilde{x}\|$
$\kappa$ -2.0 -1.5 -1.0 -0.5 0.5
$\mu=0.1$ 0.86 0.86 0.86 0.86 0.86
$\mu=0.3$ 2.84 2.84 2.84 2.84 2.84
$\mu=0.5$ 5.26 5.26 5.26 5.26 5.26
$\mu=0.7$ 8.30 8.30 8.30 8.30 8.30
$\mu=0.9$ 12.22 12.22 12.22 12.22 12.22